OpenStudy (mikezack123):

Which graph represents the solution set for the inequality 16 ≥ -3x + 4 > -17?

OpenStudy (texaschic101):

16 >= -3x + 4 > -17 -- subtract 4 16 - 4 >= -3x + 4 - 4 > -17 - 4 12 >= -3x > -21 -- divide by -3 and change the sign 12/-3 <= -3x/-3 < -21/-3 -4 <= x < 7 Is this a line graph ?

OpenStudy (texaschic101):

if so, then we have x greater then or equal to -4, and less then 7 closed circle on -4, open circle on 7, shading in between

OpenStudy (mikezack123):

@texaschic101

OpenStudy (mikezack123):

yep, its a line graph. Thank you very much! :)